1. You want to determine the linear order and genetic distance between three genes on chromosome ll in Drosophila. The recessive mutant alleles are g, m, and o. You cross heterozygous females to homozygous recessive males, score the phenotypes of progeny and obtain the following: 540 g m 560 O 162 m 170 g o 245 g 240 m o 46 g m o 37

a) Determine the orientation of mutant vs. wild type alleles on the female parent’s chromosomes. (1pt)

b) Determine gene order (remember, what matters is which gene is in the middle). (1pt)

c) Calculate the distance between the three genes and draw the genetic map. (1pt)

d) Calculate interference and provide a brief verbal interpretation of your result. (1pt)
Answer #1

Answer:

genotypes

Number of progeny

g m +

540

+ + o

560

+ m +

162

g + o

170

g + +

245

+ m o

240

g m o

40

+ + +

37

Total

2000

Parental combinations are more so, the parental combinations are g m + & + + o

1. If single cross over (SCO) occurs between g & m

Normal order= g ---------m & + ----- +

After cross over= g ----- + & +------m

g ----- + recombinants are 170+245= 415

+ ------m recombinants are 162+240=402

Total recombinants = 817

RF = (817/2000.)*100 =40.85%

2. If single cross over (SCO) occurs between m & +

Normal order= m --------- + & + --------- o

After cross over= m --------- o & + ------ +

m --------- o recombinants are 240+46= 286

+ ------ + recombinants are 245+37= 282

Total recombinants = 568

RF = (568/2000)*100 = 28.4%

3. If single cross over (SCO) occurs between g & +

Normal order= g --------- + & +------o

After cross over= g --------- o & + --------- +

g --------- o recombinants are 170+46=216

+ --------- + recombinants are 162+37=199

Total recombinants = 411

RF = (411/2000)*100 = 20.55%

% RF = Map unit distance

The order of gene is -----

g ---------20.55 m.u.---------- o -----------28.4 m.u.----------m

Expected double cross over frequency (EDCO) = RF b/e g & o * RF b/e o & m

= 0.205 * 0.284= 0.058

Observed double cross over frequency (ODCO)= (46+37)/2000 =0.042

Coefficient of Coincidence (coc)= EDCO/ODCO = 0.042/0.058=0.72

Interference= 1-coc= 1-0.72=0.28

Interference is positive

Female heterozygous genotype= g + m/+ o +